# Multi-orbital two-particle self-consistent approach – strengths and limitations

### v1

@@ Appendix B Comparision to ED
In this appendix, we present further comparisons to ED. On the one hand, we show data at lower temperatures, which demonstrate that the results improve since the Hartree-Fock approximation becomes more reliable, see Fig. 7. Furthermore, by reducing the filling we also find better agreement with ED in the dimer case, see Fig. 6a, while adding an inter-orbital coupling at half filling does not improve the results, see Fig. 6b.
Figure 6: Comparison between TPSC and ED at $\beta=2$, for different $U$ at $J=0$ at quarter filling (panel a) and with an inter-orbital coupling of unit strength (panel b). For these comparisons we include the zeroth order self-energy correction.
@@
Figure 7: Comparison between TPSC and ED at $\beta=2$, for different $U$ and $J$ combinations. For these comparisons we include the zeroth order self-energy correction.

### v3

@@ Appendix E Comparison to ED
In this appendix, we compare the results from TPSC with ED. Due to the restriction to finite-size systems in ED, this is essentially a comparison of Hartree-Fock with ED, since the fixing of the local double occupancies only gives minor corrections on top of Hartree-Fock. This comparison is complementary to the analytical considerations in Sec. 3.1. There, we traced back the breakdown of TPSC to the incorrect description of the local thermal expectation values by Hartree-Fock. To emphasize this, we show data at two temperatures $\beta=0.5t,2t$. These results demonstrate that the results improve at lower temperatures since the Hartree-Fock approximation becomes more reliable, see Fig. 6. Furthermore, by reducing the filling we also find better agreement with ED in the dimer case, see Fig. 7a, while adding a kinetic inter-orbital coupling at half filling does not improve the results, see Fig. 7b.
The model we study is again akin to Fig. 1 and its Hamiltonian is given by
where we use the same Hubbard-Kanamori interaction as in the main text, see Eq. (47). As in the main text $U$ and $J$ are measured in units of $t$, which is set to $1$. We consider the weak coupling limit in which TPSC (Hartree-Fock) should give the correct leading order behavior. For the ED calculations, we use pyED [113] from the TRIQS framework [114]. The results for all three methods as a function of $U$ and $J$ are plotted in Fig. 6.
@@
Figure 6: Comparison between TPSC3 (blue), TPSC5 (yellow) and ED (red) for the model defined in Eq. (144) at $\beta=0.5$ on the left and $\beta=2$ on the right, for different $U$ and $J$ combinations. We include the zeroth order self-energy in both TPSC variants.
Both TPSC3 and TPSC5 show indeed the correct leading order behavior for all TPEVs, irrespective of the ratio of $U$ and $J$. Furthermore, we observe that TPSC5 performs significantly worse when $J$ becomes unphysically large, while TPSC3 stays much closer to the ED result. The TPEV do match well between TPSC3 and ED throughout the whole range of examined parameters. In general, the deviation is strongest in the pair-hopping component, in which the error monotonously increases with $J$. We stress again that since we stay in the weak-coupling regime, these benchmarks have no direct implications for applications in the strongly correlated regime.
To complement the results presented in Fig. 2, we show the same comparison of ED and the TPSC variants at lower filling (upper panel) and with additional inter-orbital kinetic coupling introduced in Fig. 7. The results do qualitatively agree with Fig. 2. As expected, the TPEV are closer to ED at lower filling.
Figure 7: Comparison between TPSC and ED at $\beta=2$, for different $U$ at $J=0$ at quarter filling (panel a) and with a kinetic inter-orbital coupling of unit strength (panel b). For these comparisons we include the zeroth order self-energy correction.